OpenStudy (anonymous):

Which of the following is the slope between the points (0, -7) and (4, 12)

hero (hero):

Hint: Use y = mx + b

hero (hero):

-7 = 0m + b 12 = 4m + b Solve the first equation for b to get -7 = b Substitute b = -7 into the second equation to get 12 = 4m - 7 Add 7 to both sides to get 12 + 7 = 4m 19 = 4m Divide both sides by 4 to get 19/4 =m or 4.75 = m